MSVF operational status:
IDA engineers visited the station to decommission the KS54000 and replace it with an STS-6 borehole seismometer. An infrasound sensor, QEP and WXT weather station were also installed.
The station operator installed two upgraded Q330s.
The station is back online. Details regarding power restoration not yet known.
The station is down due to power outages caused by tropical cyclone Ana. When possible, station operators will visit the station to assess damages and restore power.
The station is back online as of 2019,066-08:47 after station operators cycled 48v DC power to the vault from the recording room aux box.
The station is down as of day 2019,064-02:26 due to communication problems between the digitizers and station computers. IDA engineers are working with the station operator to resolve this issue.
The station is back up as of day 2018,060-11:10 after an IDA engineer visited the station and replaced the LAN extenders. The msvf00 digitizer was also replaced on day 2018,066 (Mar. 7).
The station operator visited the station and reports faulty LAN extenders between the digitizers and station computers, and that the Newmar was replaced on January 12. IDA engineers are ordering replacement LAN extenders.
The station is down as of day 2017,329-14:31 due to a power supply failure. IDA engineers are working with the station operators to replace the Newmar power supply.
The station was down from day 2017,079-04:44 to 2017,081-05:21 due to a power outage. Data acquisition was restored after the station operator visited the station and reset the circuit breaker.
The station is back online as of day 2017,045-22:49 after the station operator replaced the Newmar power supply.
The station is down as of day 2017,042-04:06 due to power problems at the station. A faulty power supply is suspected.
The station wa down from day 2016,080-22:31 to 090-00:00 due to power problems following the passage of Cyclone Winston over the island. Data acquisition was restored after the station operator visited the station and reset the circuit breaker.
An IDA engineer visited the station to upgrade the power system and replace the station computers.
IDA engineers are onsite to install a Trillium120 posthole
seismometer, a strong motion sensor, and Paroscientific
IDA station MSVF is back online as of day 2013,277-04:45 after
IDA engineeers visited the station and replaced the DAS.
An IDA engineer visited the station and solved the power problem. However, there are still GPS problems, and communication problems with the digitizers. The station is down pending another site visit by an IDA engineer.
IDA station MSVF is currently down due to a power outage. Station operator reports problem with UPS.
IDA station MSVF has been upgraded to a Quanterra Q330HR system. Sensors installed include KS54000 and Trillium 120 posthole seimometers, and an Episensor strong motion sensor.
IDA station MSVF is down since day 2012,158-12:41 due to communication problems between the DAS and the ISP. IDA engineers are investigating.
An IDA engineer is on-site and has restored power and telemetry to the station. Data is seismic as of day 2012,069-22:28.
IDA station MSVF is down as of day 2012,005-05:04 due to power problems at the station. IDA engineers are working with the station operator to resolve the power problems.
The station is back on-line after the station operator replaced the GPS clock with one shipped from UCSD.
The station operator reports cycling power to the GPS. However, the GPS clock remains unlocked, and there are problems with the DAS yet to be determined.
The station is down again due to problems with the GPS clock. The station operator is requested to cycle power to the GPS clock.
The station is back on-line as of day 2010,355-02:51 after the GPS clock became locked again.
The station is down due to problems with the GPS clock. The station operator will be requested to cycle power to the GPS clock.
An IDA engineer visited the station and repaired the power system, and replaced the seismometer interface box and reconnected the strong motion sensor.
The station is down since day 2010,098-06:42 due to a power outage. A faulty circuit breaker in the recording room is suspected.
The station was down from day 2010,089-08:23 to 2010,090-21:52 due to a power problem.
The station is back online as of day 2010,033-02:30. The station was down due to a power problem.
The station is down as of day 2010,021-04:09 for reasons yet to be determined.
The station is back on-line as of day 2010,010-23:59. The GPS clock is locked.
The station is down as of day 2009,348-05:10 due to power problems. The GPS clock is also in an unlocked state.
The station is back on-line as of day 2009,341 after the station operator restored power to the station.
The station is down as of day 2009,335 due to power problems that are yet to be determined.
The triggered channels from the Episensor strong motion sensor are non-seismic for reasons yet to be determined.
The station is online again after an IDA engineer visited the station and fixed problems with the power system. The broadband seismometer data are seismic again.
An IDA engineer visited the station and restored data acquisition and established a telemetry link to the station. However, the station remains down due to problems with the KS-54000 seismometer.
The station remains down. An IDA engineer is on-site and investigating a problem with the uplink cables and the digitizer.
The station is down as of about day 190-00:59 due to communication problems between the DAS and the ISP that are yet to be determined.
The station is up after an IDA engineer restored power and upgraded the station to a MK8 system. Equipment including a new ISP, power supply, and microbarometer were installed.
Station is down due to problems with the Main Power Supply.
Station was down for time period yet to be determined due to power surge which damaged the remote power supply. Station operable again after an IDA engineer fixed the power supply, replaced circuit boards and DAT drive.
Station is back up after station operator installed new ARS and DAT drive on 2003.11.04.
Station is down due to problems with the datalogger. The ARS is not taking tapes.
Station up, data recording restored after clock replaced.
Another replacement power supply installed. Problems with external clock.
Replacement power supply delivered 2003.01.29, installed and immediately failed.
The station operator reports the main power supply has failed. A spare power supply is being prepared for shipment to Suva for installation by the FEA.
The RT303 card recording data from the KS54000 vertical channel is behaving anomalously.
Station operation is restored after IDA engineers installed a new wellhead cover and re-installed the data acquisition system.
New EPROMs installed to correct DSP problem.
Station down since 2002.02.24 for rebuilding of the wellhead pad. KS-54000 vertical channels now behaving normally. A field trip to re-install the acquisition system will take place shortly.